OpenStudy (anonymous):

HELP I WILL FAN AND MEDAL Find the approximate area of a circle with a diameter of 14 meters. use 22/7 for pi A. 62 square meters B. 154 square meters C. 44 square meters D. 591 square meters

OpenStudy (anonymous):

@kinjalojha

OpenStudy (anonymous):

radius r = 14/2 = 7 m area = pi r^2= 22/7 x 49 = 154 sq m..

OpenStudy (anonymous):

ans is B

OpenStudy (anonymous):

ok ok so is it ok if you help me with all of the question i need help with cause your really good

OpenStudy (anonymous):

ive already faned you so here is your meal

OpenStudy (anonymous):

thanx... :)

OpenStudy (anonymous):

if u need help.. ask me ..